OpenStudy (anonymous):

Use part I of the Fundamental Theorem of Calculus to find the derivative of y=∫sqrt(x) to -6 cos(t)/(t^7)dt

OpenStudy (anonymous):

sqrt(x) is on the top and -6 on the bottom. sorry.

OpenStudy (perl):

ok like this $$ \Large { \int_{6}^{\sqrt{x}} \frac{\cos(t)}{t^7}~dt } $$

OpenStudy (anonymous):

yeah

OpenStudy (perl):

we can use the fundamental theorem of calculus $$ \Large \rm { \frac{d}{dx}\int_{a}^{x} f(t) dt = f(x) }$$

OpenStudy (perl):

but theres a slight hitch, the upper limit is sqrt(x), not x

OpenStudy (perl):

we have to make a chain rule modification

OpenStudy (anonymous):

oh ok. i put in cos(sqrt(x))/sqrtx^(7) but the computer marked me wrong

OpenStudy (perl):

Fundamental theorem of calculus take two, (with chain rule) $$ \Large \rm { \frac{d}{dx}\int_{a}^{u(x)} f(t) dt = f(u(x)) \cdot u '(x) }$$

OpenStudy (anonymous):

did you mean f'(u(x))*u'(x) ?

OpenStudy (perl):

$$ \Large { \frac{d}{dx}\int_{6}^{\sqrt{x}} \frac{\cos(t)}{t^7}~dt = \frac{\cos(\sqrt{x})}{(\sqrt{x})^7} \cdot \frac{1}{2\sqrt{x}} } $$

OpenStudy (perl):

i meant f(u(x)) * u'(x)

OpenStudy (perl):

here is an explanation with some examples http://www.sosmath.com/calculus/integ/integ03/integ03.html

OpenStudy (anonymous):

hmm for some reason they're marking it wrong. Did i plug it in wrong?

OpenStudy (perl):

yes the square root x should be in the denominator

OpenStudy (anonymous):

ok i see thank you!!

OpenStudy (perl):

$$ \Large { \frac{d}{dx}\int_{6}^{\sqrt{x}} \frac{\cos(t)}{t^7}~dt = \frac{\cos(\sqrt{x})}{(\sqrt{x})^7} \cdot \color{blue}{\frac{1}{2\sqrt{x}}} } $$

OpenStudy (anonymous):

yes that's the right answer